OpenStudy (anonymous):

Find the pH of: an aqueous solution containing 0.073 g HCl in 2.00 L water

OpenStudy (anonymous):

given mass = 0.073 g find moles of HCL = 0.073 / 36.5 = 0.002 moles find molarity by dividing 0.002 by 2 molarity = 0.001 Ph = -log (H+) ph = -log(0.001) solve this to find the Ph

OpenStudy (anonymous):

ok, thanks for the working but why can you use log (0.001)? like why does (H+) = molarity of 0.001?

OpenStudy (anonymous):

molarity gives the concentration of a solution ...